### v3

@@ 1.2.3 Capabilities of 6G
Regarding development goals, 6G aims to achieve seven significant objectives: inclusivity, ubiquitous connectivity, sustainability, innovation, security/privacy/resilience, standardization and interoperability, and accessibility. It may become a new digital infrastructure that better connects the physical and virtual worlds, supports new users, and empowers new applications. Regarding performance indicators, the Recommendation specifies 15 key capability metrics for 6G, divided into two categories. The first category focuses on enhanced capabilities for IMT-2020, including peak rate, user experience rate, spectrum efficiency, regional traffic density, connection density, mobility, latency, reliability, and security/privacy/resilience performance—totaling nine indicators. The second category supports new functionalities for extending IMT-2030 use cases, encompassing coverage, perception, AI, sustainability, interoperability, and positioning—totaling six indicators. Each capability may exhibit varying relevance and applicability across different usage scenarios. The ITU’s vision for 6G encompasses a range of services and scenarios to advance communication capabilities far beyond what the current 5G technology can achieve.
Among the newly added capabilities mentioned above, the applicable AI-related capabilities draw the most attention. The possession of AI-related capabilities by 6G networks can be understood from several aspects. Firstly, it implies that the network can optimize its own performance through AI. Secondly, it indicates that the network is capable of providing support for the operation of AI. The ultimate form will be that the network can offer AI services to users and equipment just as it provides communication connections. Such capabilities can only be achieved through the in-depth fusion of wireless communication and AI.
@@ 1.3 Overview of AI and Communication Integration in 6G
In the vision and new capabilities of 6G network, the integration of AI and wireless communication is emphasized repeatedly. In this context, this section will elaborate on why and how to integrate AI and wireless communication. Generally speaking, AI will endow wireless networks with “autonomy” and “intelligence”, while wireless networks will provide AI with a broad application space, forming a two-way cyclic enabling. These two inseparable strategic development fields will continuously promote the convergent innovation and technological innovation of information, communication, and computing, and realize people’s visionary prospects in the spatiotemporal domain, information interaction types, and cross-industry convergent and innovative applications.
@@ 1.3.1 Drivers of Integrating AI into Wireless Communication
Future wireless network systems will evolve into integrated wireless infrastructure platforms deeply integrating “communication,” “perception,” “computing,” “intelligence,” and “storage.” These platforms will be capable of providing customized or personalized services on demand, surpassing the capabilities of conventional communication systems that rely on fixed architectures and predefined rules. However, such systems will need to leverage AI technologies. They will utilize big data and knowledge bases for inference and decision-making, employing models’ generalization abilities to adapt to various environments and scenarios and providing optimal resource allocation and management solutions. This section will analyze the driving forces of the integration of wireless communication and AI from the following two perspectives.
1) The diverse capabilities of AI are expected to propel the performance limits of wireless communication systems to new heights. AI possesses remarkable capabilities in handling big data. Mobile communication networks generate vast and diverse data every moment[21, 22]. The international data corporation has predicted that by 2025, the global daily data generation will reach 491 exabytes [23]. This data typically includes terminal, wireless air interface, network, and service data in various formats, such as text, images, extensible markup language (XML), hyperText markup language (HTML), graphics, and audio/video information, as shown in Table 1. Network operators can use AI technologies to train and make inference decisions on different types of mobile big data, enhancing network performance and optimizing across various dimensions and objectives.
Table 1: Data Information Table
@@
Business data Internet protocol information, uniform resource locator information, subscriber tariff data, internet access time Business identification and awareness, business anomaly detection, traffic management
@@
Besides, AI has predictive capabilities that are absent in conventional algorithms. AI can learn from historical data via ML algorithms and construct models to forecast future network requirements and possible issues. In wireless communication, this indicates the capacity to predict network congestion, device failures, or alterations in user behavior in advance. For instance, AI can anticipate potential traffic peaks during specific holidays or large-scale events and take resource allocation measures in advance, such as increasing the capacity of temporary BS or optimizing routing strategies, thereby effectively averting network breakdowns and enhancing user experience.
@@
What’s more, the adaptive capacity is also a prominent characteristic of AI. The wireless communication environment is highly dynamic and is affected by various factors such as weather, terrain, buildings, and changes in the location of mobile devices. The adaptive ability of AI enables it to make rapid adjustments according to environmental changes and fluctuations in network conditions. For example, when a user moves from indoors to outdoors or is in a high-speed moving vehicle, the signal strength and interference situation will change. AI can sense these changes in real time and automatically adjust communication parameters to ensure the stability of communication quality, adapt to different communication scenarios, and improve the reliability and flexibility of the wireless communication system.
@@
From another perspective, in contrast to the hierarchical management of different levels in conventional communication systems, AI can theoretically learn hidden structures and parameters to fit arbitrarily complex functions. This provides a more effective way to sense the complex and variable wireless environment and characterize the network state space. On the other hand, the designs of different communication system modules might have conflicting goals, and there might be performance constraints between modules. There are trade-offs between performance metrics, such as channel capacity and interference, transmission reliability, and system energy consumption. Optimizing each module individually often fails to achieve overall optimal performance. In such cases, AI can facilitate joint optimization design among modules. By leveraging advanced AI technologies like DL, different modules can learn and optimize together, thus achieving a more optimized overall system performance.
2) AI presents a novel and promising avenue for resolving the aforementioned challenges encountered in wireless communications. The core reason for the challenges summarized in Section 1.1 of this paper is that the current wireless network are not intelligent enough. Future wireless network systems will evolve into an integrated wireless infrastructure platform that deeply integrates “communication”, “sensing”, “computation”, “intelligence”, and “storage”. It will provide personalized services, which are far beyond the processing and management capabilities of conventional artificial rule pre-definition and execution. Therefore, communication systems require an automated and intelligent architecture and process to ensure the operation and development of networks and services. AI technology, which makes inferences and judgments based on big data and knowledge bases and utilizes the generalization ability of models to adapt to various environments and scenarios, can provide optimal resource allocation and management solutions and is the most promising way to achieve the above goals.
@@
Specifically, on the wireless radio access network (RAN) side, AI can be combined to conduct data analysis and decision-making for the control unit plane of 5G new radio (NR), realizing the immediate dynamic performance optimization of wireless resource scheduling and management. The next generation node B (gNB) BS can achieve intelligent on/off switching based on AI to achieve dynamic energy savings. On the CN side, with the network data analysis function (NWDAF) network elements achieving comprehensive interconnection with surrounding CN elements and realizing data collection, AI will also participate in the control and decision-making of the CN, including quality of service (QoS) guarantee, traffic handing, 5G edge computing, and load balancing of network functions.
International communication standard organizations, such as ITU, 3rd generation partnership project (3GPP), and IMT-2030, have initiated the standardization of AI applications in communication systems, and significant progress has been made. In the future, AI will further empower the infrastructure within networks, network management and operation systems, and service systems, fully unleashing the potential of integrating wireless communication and AI in the communication ecosystem and applications.
@@ 1.3.2 Vision of AI and 6G Integration
The advancement of AI technology has significantly empowered the development of wireless networks. The application of AI to enhance wireless networks is generally referred to as AI for Network (AI4NET). Its essence lies in using AI to improve wireless networks’ performance, efficiency, and user experience. Research in AI4NET primarily focuses on utilizing AI to optimize conventional algorithms, enhance network functionalities, and improve network operations and management. Figure 3 illustrates the degree of integration between 6G and AI from three perspectives: AI4NET, network for AI (NET4AI), and AI as a service (AIaaS). AI4NET’s application in 5G networks has significantly facilitated the intelligent development of mobile communication networks and vertical industries. Its role primarily manifests in operational and maintenance intelligence and network element intelligence [24].
The former emphasizes utilizing AI to optimize conventional algorithms and automate and intellectualize tasks across network layers and operations. For instance, AI empowers wireless air interface transmission to achieve adaptive channel state information (CSI) feedback, beam management, and other enhancements to transmission performance. The latter focuses on proposing AI-driven intelligence for individual network elements and functional entities, breaking down proactive network responses into predictions of multiple performance metrics. This approach ultimately aims to optimize network resource allocation and reduce network latency. Examples include AI-based traffic forecasting, mobility management, and anomaly detection applications. However, in 5G networks, AI4NET mainly operates in an “add-on” mode, introducing additional latency and data processing bottlenecks. This approach somewhat hinders AI applications’ effectiveness and can raise concerns regarding data privacy and security issues.
@@
AI’s evolution and developmental needs in wireless networks point towards 6G networks no longer serving as communication networks. Instead, they will transform into integrated information networks that combine communication, perception, and computing into intelligent services. 6G networks will further deepen the integration of AI with wireless networks. While AI empowers 6G networks, 6G networks will also empower AI, achieving NET4AI, which provides comprehensive support for the deployment and application of AI in wireless environments [25]. The essence of NET4AI is to provide various support capabilities through the NET4AI, enabling more efficient and real-time AI training/inference and enhancing data security and privacy protection. An illustration of AI-native 6G has been shown in [26], which presents a vision of a new programmable network device architecture for the NET4AI phase. The authors of [27] proposed a novel network architecture for providing native support for AI, called an AI-oriented network, which is represented as a network management framework with distributed AI computational capabilities and multi-party participation built in 6G networks. The authors of [28] combined two emerging key technologies, DT and generative AI, to provide a new way of thinking about the intelligence of 6G networks. The authors of [25] briefly described the two phases of AI4NET and NET4AI at the link and network levels and argued that NET4AI will be the backbone of 6G and future wireless networks.
Moreover, 6G networks are expected to incorporate an AI plane from its inception and other newly added logical planes such as computing, data, and security. These will be interfaced in the architecture design to support functionalities like AI plane connectivity, AI plane data operations, AI plane computing capabilities, and AI plane security features. This expansion moves beyond conventional network scopes of connectivity services to encompass capabilities in computing power, data handling, algorithms, and more. The goal is to fully embed the intelligence provided by AI into the interconnection functionality of 6G networks. Additionally, through the data flow capabilities provided by 6G networks and various vertical tasks, constructing and training new AI models’ architecture NET4AI aims to achieve native AI.
@@
Furthermore, to effectively support “inherent intelligence” and achieve ubiquitous “Native AI,” 6G networks will treat AIaaS for provision and processing. This concept gives rise to the idea of 6G AIaaS. Specifically, 6G AIaaS utilizes resources and functionalities within the network (including 6G CNs, wireless access networks, and terminals), such as connectivity, computing, data, and models. It aims to construct a distributed, efficient, energy-efficient, and secure AI service ecosystem. This ecosystem includes AI model training, inference, deployment, and other functionalities in a low-carbon open environment. The transition from network-oriented AI in 5G to AI-oriented 6G networks aims to provide network experiences that surpass data-centric approaches and directly cater to user needs [29]. Not only can it redefine the ecosystem of edge cloud, but it can also build new business models through 6G mobile networks, transitioning from past connectivity-oriented networks to service-oriented networks, ultimately achieving ubiquitous intelligence [30, 31].
AIaaS’s typical scenarios include but are not limited to, smart cities, smart agriculture, universal education, and smart industry. Typical applications include unmanned taxi services, smart grid inspections, home health monitoring, and virtual classrooms. The specific AI services involved can be categorized from different perspectives. On the one hand, from the perspective of object orientation, they can be divided into two main categories. One category provides AI services for 6G networks, which are used to optimize network performance, corresponding to AI4NET mentioned earlier. Examples include using end-to-end AI for customized air interface and network optimization and automated maintenance, leveraging AI trained on historical user data to predict traffic and optimize network resources. The other category provides AI services for end-users and third-party users (e.g., hospitals, autonomous vehicles, factories), corresponding to NET4AI, as previously discussed. For example, in industrial internet scenarios, wireless networks provide AIaaS to automatically identify industrial product quality using AI image-related technologies and trigger timely actions by industrial robots using reinforcement learning (RL) [32] technologies. Each industrial detection scenario may require fine-tuning ML algorithms with a long deployment cycle. 6G networks enable rapidly deploying targeted AI services through large-scale distributed training, integration, and inference. On the other hand, AI services come in various types. According to the different stages of the AI model’s lifecycle, they can typically be classified into data-oriented AI services, training-oriented AI services, inference-oriented AI services, and validation-oriented AI services. These AI services vary in emphasis, quality assessment metrics, and key performance indicators (KPI).
The quality of AI service (QoAIS) is a multidimensional evaluation and assurance framework within AIaaS, quantifying and grading user demands for AI service quality at the user level. QoAIS integrates into the intelligent native network of 6G to guide the provision of corresponding AIaaS and ensure user requirements are met. However, QoAIS cannot directly adopt conventional QoS metrics from communication networks, which primarily focus on performance aspects of connectivity services such as transmission rates and latency. With the evolution of 6G networks, alongside conventional communication resources, multidimensional resources will be integrated, including storage, distributed heterogeneous computing capabilities, data, and AI algorithms. Therefore, assessing the quality of native AI services within the network requires consideration across multiple dimensions, including connectivity, computing capabilities, algorithms, and data. As concerns over data privacy protection and network autonomy grow within the industry, the design of the QoAIS metric system should also incorporate dimensions such as security, privacy protection, and autonomous management [33].
@@
Moreover, with the continuous advancement of AI technology, large language models (LLM) [34] represent a significant breakthrough in the field of NLP, contributing to the development of 6G AI. Compared to conventional small-parameter models, LLM demonstrates robust contextual understanding, coherent text generation, logical reasoning, and generalization capabilities. Existing universal large models, such as GPT, can handle various domain-specific problems, demonstrating competence in addressing knowledge and issues within the communication domain as well [35]. The emergence of LLM can bring a new paradigm to integrating 6G and AI. The intrinsic AI architecture in 6G can provide linkage, computation, model decomposition, and model distribution services for the training and inference of LLM. LLM can empower various domains of 6G networks (e.g., air interface, network side, network security, network O&M, etc.), enabling intelligent services and closed-loop control. In 6G networks, customized specialized large models can be deployed across different network layers. These specialized large models can collaboratively address network issues through task decomposition and composition, cross-layer collaboration, cloud-edge collaboration, and other modes, thereby enhancing the intelligence level of the network.
In addition to using customized AI models tailored for specific functionalities and tasks, large models demonstrate superior generalization performance across multiple tasks and breakthrough capabilities on complex tasks that small models cannot achieve. The functionalities of large models increasingly align with the multi-scenario, multi-task characteristics of 6G, effectively alleviating the workload in model perception and performance metric setting within 6G networks and showing broad application prospects. A possible form of integrating 6G with AI is illustrated in Figure 4. By leveraging data from diverse heterogeneous networks within 6G, these large models can synthesize different 6G scenarios and services, laying a crucial foundation for AI empowerment in 6G networks. For example, the proposal of 6GANA in the 6GNETGPT white paper has built an LLM similar to ChatGPT, which provides a new paradigm for advanced operations and management of 6G networks [31].

### v4

@@ 1.2.3 Capabilities of 6G
Regarding development goals, 6G aims to achieve seven significant objectives: inclusivity, ubiquitous connectivity, sustainability, innovation, security/privacy/resilience, standardization and interoperability, and accessibility. It may become a new digital infrastructure that better connects the physical and virtual worlds, supports new users, and empowers new applications. Regarding performance indicators, the Recommendation [21] specifies 15 key capability metrics for 6G, divided into two categories. The first category focuses on enhanced capabilities for IMT-2020, including peak rate, user experience rate, spectrum efficiency, regional traffic density, connection density, mobility, latency, reliability, and security/privacy/resilience performance—totaling nine indicators. The second category supports new functionalities for extending IMT-2030 use cases, encompassing coverage, perception, AI, sustainability, interoperability, and positioning—totaling six indicators. Each capability may exhibit varying relevance and applicability across different usage scenarios. The ITU’s vision for 6G encompasses a range of services and scenarios to advance communication capabilities beyond what the current 5G achieves.
Among the above-mentioned newly added capabilities, the applicable AI-related capabilities have drawn the most attention. The possession of AI-related capabilities by 6G networks can be understood from several aspects. Firstly, the network can optimize its own performance through AI. Secondly, the network is capable of providing support for the operation of AI. Eventually, the network can offer AI services to users and equipment just as it provides communication connections. Such capabilities can only be achieved through the in-depth fusion of wireless communication and AI.
@@ 1.3 Overview of AI and Communication Integration in 6G
This section elaborates on why and how to integrate AI and wireless communication. Generally speaking, AI can endow wireless networks with “autonomy” and “intelligence”, while wireless networks provide AI with a broad application space. These two inseparable strategic development fields will continuously promote the convergent innovation and technological innovation of information, communication, and computing, and realize people’s visionary prospects in the spatiotemporal domain, information interaction types, and cross-industry convergent and innovative applications.
@@ 1.3.1 Drivers of Integrating AI into Wireless Communication
Future wireless network systems will evolve into integrated wireless infrastructure platforms deeply integrating “communication”, “perception”, “computing”, “intelligence”, and “storage.” These platforms will be capable of providing customized or personalized services on demand, surpassing the capabilities of conventional communication systems that rely on fixed architectures and predefined rules. Such systems will need to leverage AI technologies. They will utilize big data and knowledge bases for inference and decision-making, employing models’ generalization abilities to adapt to various environments and scenarios and providing optimal resource allocation and management solutions. This section analyzes the driving forces of the integration of wireless communication and AI from the following two perspectives.
Table 1: Data information
@@
Business data IP information, uniform resource locator information, subscriber tariff data, internet access time Business identification and awareness, business anomaly detection, traffic management
(1) AI to push limits of wireless communication systems. AI possesses remarkable capabilities in handling big data. Mobile communication networks generate vast and diverse data every moment[24, 25]. The international data corporation has predicted that by 2025, the global daily data generation will reach 491 exabytes [26]. These data typically include terminal, wireless air interface, network, and service data in various formats, such as text, images, extensible markup language, hyperText markup language, graphics, and audio/video information, as shown in Table 1. Network operators can use AI to train and make inference decisions on different types of mobile big data, enhancing network performance and optimizing across various dimensions and objectives.
@@
AI has predictive capabilities absent in conventional algorithms. AI can learn from historical data via ML algorithms and construct models to forecast future network requirements and possible issues, including network congestion, device failures, or alterations in user behavior. For instance, AI can anticipate potential traffic peaks during specific holidays or large-scale events and take resource allocation measures in advance, such as increasing the capacity of temporary BS or optimizing routing strategies, thereby effectively averting network breakdowns and enhancing user experience.
@@
Adaptivity is a prominent characteristic of AI. The wireless communication environment is highly dynamic and affected by various factors, such as weather, terrain, buildings, and changes in the location of mobile devices. The adaptive ability of AI enables it to make rapid adjustments according to environmental changes and fluctuations in network conditions. For example, when a user moves from indoors to outdoors or is in a high-speed vehicle, the signal strength and interference situation will change. AI can sense these changes in real time and automatically adjust communication parameters to ensure the stability of communication quality, adapt to different communication scenarios, and improve the reliability and flexibility of wireless communication systems.
@@
In contrast to the hierarchical management of different levels in conventional communication systems, in theory, AI can learn hidden structures and parameters to fit arbitrarily complex functions. This provides a more effective way to sense the complex and variable wireless environment and characterize the network state space. On the other hand, the designs of different communication system modules might have conflicting goals, and there might be performance constraints between modules. There are trade-offs between performance metrics, such as channel capacity and interference, transmission reliability, and system energy consumption. Optimizing each module individually often fails to achieve overall optimal performance. In such cases, AI can facilitate joint optimization design among modules.
(2) AI for resolving challenges in wireless communications. The core reason for the challenges summarized in Section 1.1 is that the current wireless network is not intelligent enough. Future wireless network systems will evolve into an integrated wireless infrastructure platform that deeply integrates “communication”, “sensing”, “computation”, “intelligence”, and “storage”.
@@
On the wireless radio access network (RAN) side, AI can be combined to conduct data analysis and decision-making for the control unit plane of 5G new radio (NR), realizing the immediate dynamic performance optimization of wireless resource scheduling and management. The next generation node B (gNB) BS can achieve intelligent on/off handoff based on AI to achieve dynamic energy savings. On the CN side, with the network data analytics function (NWDAF) network elements achieving comprehensive interconnection with surrounding CN elements and realizing data collection, AI will also participate in the control and decision-making of the CN, including quality of service (QoS) guarantee, traffic handing, 5G edge computing, and load balancing of network functions. International communication standard organizations, such as ITU, 3rd Generation Partnership Project (3GPP), and IMT-2030, have initiated the standardization of AI applications in communication systems, and made significant progress. In the future, AI is anticipated to further empower the infrastructure within networks, network management, operation systems, and service systems, fully unleashing the potential of integrating wireless communication and AI in the communication ecosystem and applications.
@@ 1.3.2 Vision of AI and 6G Integration
The advancement of AI technology has significantly empowered the development of wireless networks. Figure 3 illustrates the degree of integration between 6G and AI from three perspectives: AI for Network (AI4NET), Network for AI (NET4AI), and AI as a Service (AIaaS). The application of AI to enhance wireless networks is generally referred to as AI4NET. Its essence lies in using AI to improve wireless networks’ performance, efficiency, and user experience. AI4NET’s application in 5G networks has significantly facilitated the intelligent development of mobile communication networks and vertical industries. Its role primarily manifests in O&M intelligence and network element intelligence [27]. The former emphasizes utilizing AI to optimize conventional algorithms and automate and intellectualize tasks across network layers and operations. The latter focuses on proposing AI-driven intelligence for individual network elements and functional entities, breaking down proactive network responses into predictions of multiple performance metrics. This approach ultimately aims to optimize network resource allocation and reduce network latency.
@@
To meet the transformative and developmental demands of AI in wireless networks, 6G will no longer be merely a communication network that serves as a connectivity enabler for intelligent services. Instead, it will evolve into an integrated information network that combines communication, sensing, and computing functionalities. While AI empowers 6G networks, 6G networks will also empower AI, achieving NET4AI, which provides comprehensive support for the deployment and application of AI in wireless environments [28, 29]. The essence of NET4AI is to provide AI with various support capabilities, enabling more efficient and real-time AI training/inference and enhancing data security and privacy protection. The IMT-2030 (6G) Promotion Group in China envisioned the 6G network architecture in [30], as an open and innovative platform for information services, offering capabilities that transcend mere connectivity. These capabilities encompass computing power networks, trusted security, sensing, and data services, which are essential for the operation of AI in networks. The authors of [31] proposed a novel network architecture for providing native support for AI, called an AI-oriented network, which is represented as a network management framework with distributed AI computational capabilities and multi-party participation built in 6G networks.
@@
Moreover, to effectively support “inherent intelligence” and achieve “Native AI”, 6G networks will treat AIaaS for provision and processing. This concept gives rise to the idea of 6G AIaaS. Specifically, 6G AIaaS utilizes resources and functionalities within the network (including 6G CN, wireless access networks, and terminals), such as connectivity, computing, data, and models. It aims to construct a distributed, efficient, energy-efficient, and secure AI service ecosystem, which includes AI model training, inference, deployment, and other functionalities in a low-carbon open environment [32]. Not only can it redefine the ecosystem of edge cloud, but also build new business models through 6G mobile networks, transitioning from past connectivity-oriented networks to service-oriented networks, ultimately achieving ubiquitous intelligence [33, 34]. AIaaS’s typical scenarios include but are not limited to, smart cities, smart agriculture, universal education, and smart industry, etc. Typical applications include unmanned taxi services, smart grid inspections, home health monitoring, and virtual classrooms.
@@
Large language models (LLMs) [35] have represented a significant breakthrough in NLP and can potentially contribute to the development of 6G AI. Compared to traditional smaller parameter models, LLMs exhibit strong contextual understanding, coherent text generation, logical reasoning, and generalization capabilities. Existing general-purpose large models, such as the closed-source GPT model developed by OpenAI and the open-source Llama model developed by Meta (formerly Facebook), can address a wide range of domain-specific problems [36, 37]. Recently, the open-source DeepSeek LLM developed by the DeepSeek team in China has potentially become a game-changer, demonstrating the feasibility of training large-scale AI models (LAMs) at unprecedentedly low cost [38]. LLMs can potentially play a critical role in communication systems, e.g., in task processing and intelligent services.
The emergence of LLMs can bring a new paradigm for the integration of 6G and Al. The intrinsic Al architecture in 6G can provide linkage, computation, model decomposition, and model distribution services for the training and inference of LLMs. LLMs can empower various domains of 6G networks (e.g., air interface, network side, network security, network O&M, etc.), enabling intelligent services and closed-loop control. In 6G networks, customized specialized large models can be deployed across different network layers. These specialized large models can collaboratively address network issues through task decomposition and composition, cross-layer collaboration, and cloud-edge collaboration, enhancing the intelligence level of the network.
In addition to using customized AI models tailored for specific functionalities and tasks, large models demonstrate superior generalization performance across multiple tasks and breakthrough capabilities on complex tasks that small models cannot achieve. The functionalities of large models increasingly align with the multi-scenario, multi-task characteristics of 6G, effectively alleviating the workload in model perception and performance metric setting within 6G networks and showing broad application prospects. A possible form of integrating 6G with Al is illustrated in Figure 4. Leveraging data from diverse heterogeneous networks within 6G, these large models can synthesize different 6G scenarios and services, laying a crucial foundation for Al empowerment in 6G networks. For example, the proposal of 6GANA in the 6G NETGPT white paper has built an LLM similar to ChatGPT, which provides a new paradigm for advanced operations and management of 6G networks [34].